OpenStudy (anonymous):

Factor the Following Polynomials: x²-81 x²-13x+36 Which polynomial is a special product and why?

OpenStudy (anonymous):

need someone to walk me through step by step

OpenStudy (anonymous):

x^2-81=0 x^2=81 x=9

OpenStudy (anonymous):

okay so thats not the special product right?

OpenStudy (anonymous):

what do u mean by special product?

OpenStudy (anonymous):

Which polynomial is a special product and why?

OpenStudy (anonymous):

it says it in the question idk what it means

OpenStudy (anonymous):

lets do the second one, how did you start

OpenStudy (anonymous):

then mine is special

OpenStudy (anonymous):

i mean first one equation.

OpenStudy (anonymous):

\[x ^{2 } + 9^{2} \] can you do it now?

OpenStudy (anonymous):

x=9 is not the answer

OpenStudy (anonymous):

okay so I'm at x^2+9^2 now what do i do

OpenStudy (anonymous):

Oops.. it x^2-9^2 not +

OpenStudy (anonymous):

ok i am doing

OpenStudy (anonymous):

x^2-9^2=(x+9)(x-9)

OpenStudy (anonymous):

okay now what

OpenStudy (anonymous):

(x-9) and (x+9) are the factors. You've got tha answers..:D well done

OpenStudy (anonymous):

Can we move on to the second?

OpenStudy (anonymous):

\[x^2-13x+36=0\] now factorize \[x^2-9x-4x+36=0\] x and 4 common \[x(x-9)-4(x-9)=0\] bracket common \[(x-9)(x-4=0\] now x-9=0 or x-4=0 x=9 or x=4

OpenStudy (anonymous):

ok lets do the second

OpenStudy (anonymous):

2nd and 1st i have done with an easy way.

OpenStudy (anonymous):

x²-13x+36 is the question right? we don't have any common factor in the equation right now. Let's try to build.

OpenStudy (anonymous):

x²-13x+36 There is a std procedure to do such problems. Express the middle terms ( 13 ) as a sum of two no.s and last term ( 36 ) as a product. can you come up with any two no.s?

OpenStudy (anonymous):

@muhammad9t5 : You can't equate it to zero. No where it's given. Factorization and equation are two different things.

OpenStudy (anonymous):

ok